OpenStudy (anonymous):

How do you find the vertex, axis of symmetry, domain, range, x-intercept, and y-intercept of the equation: y=x^2+12x+35?

OpenStudy (anonymous):

Vertex is where the curve turns

OpenStudy (anonymous):

differentiate the function

OpenStudy (hba):

did u try to solve it

OpenStudy (anonymous):

no not yet, i don't understand how you find the vertex

OpenStudy (campbell_st):

the curve can be written in the form y = (x -h)^2 + k (h, k) is the vertex, line of symmetry is x = h \[y = (x + 6)^2 - 1\] the vertex is (-6, 1), line of symmetry x = -6 domain is all real x the range is y >= -6 y intercept let x = 0 then y = 35 x intercepts let y = 0 0 = x^2 + 12x + 35 0 = (x + 5)(x +7) the x- intercepts are x = -5, -7

OpenStudy (anonymous):

where did you get 6 from?

OpenStudy (anonymous):

@campbell_st how did you go from y = x^2 + 12x + 35 to y = (x + 6)^2 − 1 ?

OpenStudy (campbell_st):

@JoshDavoll y = x^2 + 12x + 36 -1 allows you to complete the square so its y = (x^2 + 12x + 36) -1 the original equation is y = x^2 + 12x + 35 which is factorised for the x intercepts...

OpenStudy (anonymous):

ahh thanks! :)

OpenStudy (anonymous):

thanks!! @campbell_st

OpenStudy (campbell_st):

the 6 comes from the completing the square your equation is y = x^2 + 12x + 35 by adding 1 you will get a perfect square but you also need to subtract 1 y = (x^2 + 12x + 36) - 1 is still the original equation y = (x + 6)^2 -1 is a factorised form of the original equation

OpenStudy (anonymous):

factor and you get (x+7)*(x+5) the x intercepts are thus -7, -5 the axis of symmetery is midway of that so use midpoint formula m=(x1+x2)/2 so it is (-7+(-5))/2= -6 that -6 is also the x coordinant of the vertex if you plug in -6 into (x+7)*(x+5)=(-6+7)*(-6+5)=-1 you get -1 for y coordinant of vertex so now you have the pair for the coordiant of vertex (-6, -1) to find domain you ask yourself what numbers are valid that we can plug into my function. Well it is all of them, since there is no division by zero or imaginary roots. for range you ask yourself if i had a graph where will the y values start and where will it end. x^2+12x+35 since the leading coeffecient of the highest degree term is positive it will be a "u" shaped graph. You know the vertex is (-6,-1), so it will start at -1 and go to infinity so the range is -1 to infinity.